# GPU Acceleration of SQL Analytics on Compressed Data

### v1

@@ 9. Experimental Setup and Results
We compare our tensor-based GPU query execution on compressed data against tensor-based GPU query execution on Plain data (with dictionary encoding, like TQP)  [^2] Execution plans on Plain data are an optimized version of the tensor programs generated by TQP. Plain performance are therefore equal or better than TQP., and CPU-only query execution with SQL Server and Analysis Services (Microsoft, 2025). We run GPU experiments on an Azure NC24ads A100 v4 VM (Microsoft, 2024b) having an NVIDIA A100 GPU with 80 GiB HBM, 24 vCPUs, and 220 GiB main memory. For CPU-only analytics systems, we use an Azure D64s v5 VM (Microsoft, 2024a) with 64 vCPUs and 256 GiB main memory. We report warm query times, averaged over multiple runs, after the data is loaded into memory (GPU HBM for GPU and main memory for CPU experiments) and caches have been warmed up. We monitor GPU memory usage with the nvidia-smi profiling tool.
@@
  • Else, if the RLE compression ratio $>20$, then use RLE.
  • Else, if many single-element runs exist but longer runs still yield an RLE compression ratio $>20$, then use RLE+Index.
@@ 9.1. TPC-H Queries
We evaluate our approach using a subset of TPC-H queries (Q1, Q2, Q6, Q11, Q14, Q15, Q17, Q19) at scale factors (SF) 50, 100, and 300. The TPC-H benchmark utilizes synthetic data which, compared to real-world datasets (discussed in Section 9.2), exhibits limited sparsity and redundancy. This characteristic makes techniques like V-order (Microsoft, 2024c) less effective for general compression across all queries. Therefore, to optimize for compression, we employ query-specific sort orders for the relevant tables based on the columns involved in each query’s filters and joins, as detailed in Table 7. Sorting significantly improves RLE effectiveness by increasing average run lengths and reducing the total number of runs. For instance, as illustrated for the l_returnflag column in Q1 at SF=50 (Figure 7), sorting reduces the number of runs from $>$105M to just 3, while increasing the average run length from 2.86 to slightly over 100M.
@@
Figures 8 and 9 present the peak GPU memory usage and query run times, respectively, comparing query execution performance on uncompressed (Plain) data versus compressed (RLE, Index) data formats across SF 50, 100, and 300. Our evaluation reveals substantial benefits when processing compressed data directly on the GPU. Peak memory usage (Figure 8) is significantly reduced; for example, at SF=300, Q19 requires 65.3 GiB with Plain data but only 17.5 GiB with compressed data (a 3.7×3.7$\times$3.7 × reduction). Compression enables processing larger scale factors like SF=300 for queries such as Q1 and Q17, which exceed the 80 GiB GPU memory limit with Plain data. Query run times (Figure 9) are also dramatically reduced, with speedups often exceeding 10×10$\times$10 ×. For instance, at SF=300, Q6 runs 17.3$\times$ faster (68.6 ms vs 3.96 ms) and Q19 runs 23.8$\times$ faster (470.1 ms vs 19.8 ms) on compressed data. These results highlight the effectiveness of our techniques in improving both memory efficiency and performance for TPC-H queries on GPUs.
Using SQL Server query times for SF=50 as reference, the sum of times for the 8 queries running on GPU with compressed data was 12.8x lower for SF=50 and 2.6x lower for SF=300 (6x larger SF)!
(a) Unsorted
(b) Sorted
Figure 7. Total number of runs and average run lengths (with data labels) for RLE-compressed fact table columns for TPC-H SF=50 Q1 before and after sort.
@@
Q1 l_returnflag, l_linestatus, LINEITEM
l_shipdate, l_quantity
@@
Figure 8. Peak GPU memory used while running TPC-H queries on Plain and Compressed input data for different SF.
Figure 9. TPC-H query run times, using A100 GPU, on Plain and Compressed input data for different SF.
9.2. Production queries
We evaluate our methods using representative production queries on a first-party production dataset. It has a star schema with a large fact table having 2.94 Billion rows and 13 small dimension tables. It also has 4 ‘bridge’ tables that enable joins between the fact table and 4 of the dimension tables. On disk, each table is stored as V-ordered Parquet files (Microsoft, 2024c), which is enabled by default in Microsoft Fabric. V-ordering applies sorting on tables, dictionary encoding, and other techniques to improve compression while supporting fast reads. Unlike in the TPC-H study (Section 9.1), we use the same input files, and thus the same row ordering for a given table, for all queries.
Table 8. Columns of the fact table used by the queries.
@@ 9.2. Production queries
We use three queries for this study. Table 8 shows columns of the fact table that each query uses. Q1 uses 10 columns while Q2 and Q3 use 12 columns each. Together, the queries use 15 columns which is a subset of the total columns in the fact table. Relational operations in the queries include predicate filters on dimension tables, joins and semi-joins, and group-by aggregation (SUM). Q1 includes 7 semi-joins and 2 primary-key foreign-key (PK-FK) joins, Q2 and Q3 each includes 10 semi-joins and 1 PK-FK join. Q2 and Q3 have a similar template but differ in a filter predicate.
Figure 10. Input column sizes for the 15 columns in the fact table for Plain and Compressed representations. The bottom row in the x-axis labels shows the encoding for each column in compressed representation—R: RLE, C: Composite (Plain + Index), P: Plain with bit-width reduction.

### v2

@@ 9. Experimental Setup and Results
We compare our tensor-based GPU query execution on compressed data against tensor-based GPU query execution on Plain data (with dictionary encoding, like TQP) [^2] Execution plans on Plain data are an optimized version of the tensor programs generated by TQP. Plain performance are therefore equal or better than TQP. For more details on the query optimization rules applied, see Appendix D., and CPU-only query execution with SQL Server and Analysis Services (Microsoft, 2025). We run GPU experiments on an Azure NC24ads A100 v4 VM (Microsoft, 2024b) (pay-as-you-go @\$3.673/hour) having an NVIDIA A100 GPU with 80 GiB HBM, 24 vCPUs, and 220 GiB main memory. For CPU-only analytics systems, we use an Azure D64s v5 VM (Microsoft, 2024a) (pay-as-you-go @\$3.072/hour) with 64 vCPUs and 256 GiB main memory. We report warm query times, averaged over multiple runs, after the data is loaded into memory (GPU HBM for GPU and main memory for CPU experiments) and caches have been warmed up. Our assumption is that data already resides in GPU memory during query execution. This is a key benefit of compression: data becomes smaller and can fit in GPU memory, avoiding expensive PCIe transfers. We monitor GPU memory usage with the nvidia-smi profiling tool.
@@
  • Else, if the RLE compression ratio >20>20, then use RLE.
  • Else, if many single-element runs exist but longer runs still yield an RLE compression ratio >20>20, then use RLE+Index.
@@
Figure 6. Total # of runs and average run lengths for fact table for TPC-H SF=50 Q1 before and after sort. Sorting greatly improves RLE compression, reducing runs from >>105M to 3 while increasing average length from 2.86 to >>100M.
Figure 7. Peak GPU memory usage (top) and query run times (bottom) for TPC-H queries on Plain and Compressed input data across different scale factors. Compression achieves speedups up to $23.8\times$ while reducing memory usage by up to $3.7\times$.
Figure 8. TPC-H query runtime between HeavyDB and compressed GPU execution at SF=100. Our compressed approach achieves $12.8\times$ geometric mean speedup over HeavyDB.
@@ 9.1. TPC-H Queries
We evaluate our approach using a subset of TPC-H queries (Q1, Q2, Q6, Q11, Q14, Q15, Q17, Q19) at scale factors (SF) 50, 100, and 300. The TPC-H benchmark utilizes synthetic data which, compared to real-world datasets (discussed in Section 9.2), exhibits limited sparsity and redundancy.
9.1.1. Query-Specific Data Ordering
@@
To study the impact of data ordering on RLE compression effectiveness, we employ query-specific sort orders for the relevant tables based on the columns involved in each query’s filters and joins. Each table is sorted using a global multi-column ordering (equivalent to SQL’s ORDER BY clause). The specific column orderings used for each query are detailed in Section B.1. Sorting significantly improves RLE effectiveness by increasing average run lengths and reducing the total number of runs. For instance, as illustrated for the l_returnflag column in Q1 at SF=50 (Figure 6), sorting reduces the number of runs from >105M>105M to just 3, while increasing the average run length from 2.86 to >100M>100M. In contrast, general-purpose V-order (Microsoft, 2024c) proved less effective for TPC-H even with skewness, due to high-cardinality columns (Section B.2).
@@
Figure 7 presents both peak GPU memory usage and query run times, comparing query execution performance on uncompressed (Plain) data versus compressed (RLE, Index) data formats across SF 50, 100, and 300. Our evaluation reveals substantial benefits when processing compressed data directly on the GPU. Peak memory usage is significantly reduced; for example, at SF=300, Q19 requires 65.3 GiB with Plain data but only 17.5 GiB with compressed data (a $3.7\times$ reduction). Compression enables processing larger scale factors like SF=300 for queries such as Q1 and Q17, which exceed the 80 GiB GPU memory limit with Plain data.
[…]
Using SQL Server query times for SF=50 as reference, the sum of times for the 8 queries running on GPU with compressed data was 12.8x lower for SF=50 and 2.6x lower for SF=300 (6x larger SF)! To compare with existing GPU database systems, we show HeavyDB (HeavyDB, ssed) performance in Figure 8. HeavyDB outperforms uncompressed execution on Q1 and Q6, but underperforms on the rest. Overall, HeavyDB is $12.8\times$ slower than our compressed execution (geometric mean across all queries).
9.1.2. Compression Quality Ablation Study
To study the impact of RLE compression quality on query performance, we experimented with Q17 and Q19 at SF=100. These queries demonstrate significant speedups with compression and their performance depends on the l_partkey column in LINEITEM, allowing systematic variation of compression effectiveness. We created datasets with varying compression ratios by controlling run formation. Starting with 600M LINEITEM rows naturally grouped into 20M partkeys ($\sim 30$ rows each), we systematically break runs by further dividing each row group into 2–16 smaller runs. This creates datasets with compression ratios from 30x (20M runs) to 1.87x (320M runs).
[…]
Figure 9 shows that query performance degrades as the number of runs increases and compression effectiveness decreases. With optimal compression (30x, 20M runs), Q17 executes in 28.87ms but degrades to 189.95ms with poor compression (1.87x, 320M runs)—a 6.6x slowdown. Similarly, Q19 runtime increases from 4.11ms to 24.83ms—a 6x slowdown. These results underscore the importance of good RLE compression for high query performance.
9.2. Production queries
To study the properties of production workloads, we evaluate our methods using representative production queries on a first-party production dataset, selected by the product team as a challenging case where CPU baselines perform well. It has a star schema with a large fact table having 2.94 Billion rows and 13 small dimension tables. It also has 4 ‘bridge’ tables that enable joins between the fact table and 4 of the dimension tables.Unlike in the TPC-H study (Section 9.1), we use the same input files, and thus the same row ordering for a given table, for all queries. We consider 2 ordering strategies: (1) On disk, each table has already been stored as V-ordered Parquet files (Microsoft, 2024c), which is enabled by default in Microsoft Fabric. This is not query-specific. (2) We also experiment with a query-specific sorting strategy based on cardinality ordering, sorting all query columns by cardinality (smallest first). This corresponds to real-world scenarios where BI customers optimize specific queries using simple rules.
We use three queries for this study. Q1 uses 10 columns while Q2 and Q3 use 12 columns each. Together, the queries use 15 columns (detailed column usage is provided in Table 8) which is a subset of the total columns in the fact table. Relational operations in the queries include predicate filters on dimension tables, joins and semi-joins, and group-by aggregation (SUM). Q1 includes 7 semi-joins and 2 primary-key foreign-key (PK-FK) joins, Q2 and Q3 each includes 10 semi-joins and 1 PK-FK join. Q2 and Q3 have a similar template but differ in a filter predicate.
9.2.1. Compression Effectiveness
To study compression benefits on real-world data, Figure 10 shows the sizes for the in-memory representation of the 15 columns of the fact table. The series ‘Plain’ refers to the default Plain encoding (with dictionary encoding) but with no other compression schemes applied.
Figure 10. Input column sizes for the 15 columns in the fact table for Plain and Compressed representations. The top row in the x-axis labels shows the column numbers while the bottom shows the encoding for each column in compressed representation—R: RLE, C: Composite (Plain + Index), P: Plain with bit-width reduction.
The total size for the 15 columns is 120.36 GiB with Plain and does not fit into the 80 GiB GPU HBM. Considering only the subset of columns needed by the specific queries, Q1 needs 87.53 GiB and Q2 and Q3 need 82.06 GiB, which is still larger than the GPU HBM capacity. The bottom row of the x-axis shows the encoding scheme for Compressed—‘C’: Composite (Plain + Index), ‘R’: RLE, ‘P’: Plain with bit-width reduction based on the range of values. The total size for 15 columns with Compressed is 56.84 GiB. For Q1 columns it is 43.69 GiB and for Q2 and Q3 it is 30.86 GiB, which can fit in the GPU HBM and also leave space for intermediate results. For reference, the on-disk V-ordered file size for the projected fact table for the 15 columns is 29.87 GiB.
[…]
9.2.2. Query Performance Analysis
To study performance gains on production queries, Figure 11 shows query run times on state-of-the-art CPU-based commercial RDBMS, Microsoft SQL Server, and analytical data engine, Analysis Services. Our techniques lead to substantial speedups—15.75 (Q1), 13.66 (Q2), 9.82 (Q3), 12.76 (total) over SQL Server and 8.72 (Q1), 13.16 (Q2), 8.05 (Q3), 9.52 (total) over Analysis Services. The cardinality-based sorting strategy achieved run times of 12.26 ms (Q1), 24.27 ms (Q2), and 45.19 ms (Q3), comparable to V-ordering. V-order performed better for Q2 and Q3 despite being query-agnostic, while cardinality-based sorting creates highly skewed column size distributions. This demonstrates V-order’s effectiveness without query-specific knowledge, while cardinality-based sorting provides a simple competitive alternative. In Section C.3, we present a scalability analysis showing that compressed execution enables processing datasets up to 222% (6.52B rows) of the original size within GPU memory, while Plain execution fails at 50% capacity. Compressed queries also scale better: Q2 and Q3 GPU times on Plain data at just 20% dataset size already exceed CPU times for the full 100% dataset.
9.3. Public BI Datasets
To further validate our approach on diverse real-world workloads, we evaluate our techniques using publicly available BI datasets (Vogelsgesang et al., 2018). These are collected from Tableau Public, a platform hosting over 60K user-generated BI visualizations spanning government, healthcare, transportation, e-commerce, and other domains. These datasets reflect realistic production scenarios with diverse characteristics.
Of the 34 datasets, 20 ($59\%$) contain columns meeting the compression ratio threshold (>20>20) for RLE compression. Among these 20 datasets with compression opportunities, 14 have adequate size ($\geq 1$M rows) for meaningful GPU evaluation. We evaluate the first 3 qualified queries per dataset that utilize compressible columns, resulting in 38 total queries. Unlike our TPC-H experiments which used query-specific ordering, all datasets are sorted using V-order (Microsoft, 2024c) across all columns.
Figure 12 shows compression provides significant benefits in 28 of 38 queries ($73.7\%$) with maximum speedup of $11.27\times$ and geometric mean speedup of $2.02\times$. There are also 10 queries that show slowdowns up to $3.13\times$. This occurs when queries have few RLE columns, and they are directly operated (e.g., AND operations) with plain columns; this introduces additional overhead from RLE-to-plain decompression. Overall, real-world BI queries demonstrate high compressibility and our execution framework shows effective acceleration for the majority of workloads.
Figure 12. Query runtime for public BI datasets, comparing Plain vs Compressed data execution on GPU. Compression benefits $73.7\%$ of queries with $2.02\times$ geometric mean speedup and maximum speedup of $11.27\times$.
10. Related Work
GPU acceleration for SQL analytics: There has been considerable research into using GPUs to accelerate SQL analytical queries (Shanbhag et al., 2020; Cao et al., 2023; He et al., 2022; Yuan et al., 2013; Rosenfeld et al., 2022; Yogatama et al., 2022; Sioulas et al., 2019; Yogatama et al., 2025; Yuan et al., 2025; Chrysogelos et al., 2019; Wu et al., 2025; Hu et al., 2022; Rui and Tu, 2017; Hong et al., 2025; Deng et al., 2024; Karnagel et al., 2015; HeavyDB, ssed; Data, 2025; Paul et al., 2021; BlazingSQL, 2021; Lutz et al., 2020; Maltenberger et al., 2022; Rui et al., 2020). The majority of these works, either consider input data to be already in plain format(e.g., (BlazingSQL, 2021; HeavyDB, ssed; Data, 2025; He et al., 2022)), or decompress it before executing queries (e.g.,  (Shanbhag et al., 2020)). Several of this works also only consider using GPUs to accelerate a subset of query operators (e.g., (Sioulas et al., 2019; Wu et al., 2025; Lutz et al., 2020)). Compared to this previous research, we (1) support end-to-end GPU acceleration of queries; and (2) we are the first one providing a framework allowing to run queries in GPU and directly on compressed data.
11. Conclusion
Appendix A Operator Implementation Details and Examples
This section provides detailed algorithmic implementations and concrete examples for key operations in our compressed tensor query execution system.
A.1. NOT Operator Implementation
The NOT operator requires specialized algorithms for different mask representations. For RLE masks, we compute the gaps between runs to create the inverted result (Algorithm 6). For Index masks, we convert sparse indices to continuous ranges representing all non-selected positions (Algorithm 7). These algorithms are illustrated with concrete examples below.
Algorithm 6 not_rle
Input: startstart, endend, total_size
Output: $start^{\prime}$, $end^{\prime}$
[…]
3:mask $\leftarrow$ $start^{\prime}$ $\leq$ $end^{\prime}$
4:return $start^{\prime}$[mask], $end^{\prime}$[mask]
Algorithm 7 not_index
Input: val, total_size
Output: startstart, endend
1:startstart $\leftarrow$ cat(([-1], val)) + 1
2:endend $\leftarrow$ cat((val, [total_size])) - 1
3:mask $\leftarrow$ (startstart $\leq$ endend) $\land$ (startstart << total_size) $\land$ (endend $\geq 0$)
4:return startstart[mask], endend[mask]
Figure 13. Illustration of NOT logical operator.
Example 7.
  • NOT on RLE mask. For an RLE mask with $s=[0,4]$, $e=[1,6]$, and total_size=8total\_size=8, we apply the complement_rle primitive. First, we compute the potential start points, $s^{\prime}$, for inverted ranges by concatenating [1][-1] with the original end positions and adding 1. $s^{\prime}=[0,2,7]$. Then, we compute potential end points, $e^{\prime}$, by concatenating the original start values with [total_size][total\_size] and subtracting 1. $e^{\prime}=[-1,3,7]$. We keep only the valid ranges where $s^{\prime}\leq e^{\prime}$, which gives us $s^{\prime}=[2,7]$ and $e^{\prime}=[3,7]$. This is an RLE mask covering the gaps between the original runs.
  • Step 3: NOT on Index mask. For an Index mask with $p=[2,5]$ and total_size=8total\_size=8, we apply the complement_index primitive. We compute potential start points by concatenating [1][-1] with the original indices and adding 1. $s=[0,3,6]$. We compute potential end points by concatenating the original indices with [total_size][total\_size] and subtracting 1. $e=[1,4,7]$. After applying a validity mask to ensure $s\leq e$, s<total_sizes<total\_size, and $e\geq 0$, all position ranges remain valid. This is an RLE Mask with runs covering all positions not in the original Index mask.
A.2. GroupBy Aggregation Example
We provide a detailed walkthrough of how groupby aggregation works on RLE-compressed data, illustrating the three-step process of grouping, alignment, and aggregation.
Example 8.
Consider a SQL query ”SELECT SUM(B) GROUP BY A” on RLE-compressed data as shown in Figure 14. The input consists of a group-by column A with values $[A,B,A]$ spanning runs 0–1, 2–4, 5–8 and a sum column B with value $3$, start $[0]$ and end $[8]$.
  • Step 1: Grouping. We apply the grouping operation on the group-by column to identify unique groups. Using torch.unique on the values $[A,B,A]$, we identify two unique groups (A and B) and generate an inverse index, $x=[0,1,0]$ that maps each run to its corresponding group: the first run (group-by value A) maps to group 0, the second run (group-by value B) maps to group 1, and the third run (group-by value A) again maps to group 0. This creates a temporary RLE column $g$ with $v=x$ (where $x$ is the inverse index just computed) and with $s$ and $e$ the same as that of the group-by column ($A$).
  • Step 2: Alignment. Next, we align $g$ and the aggregation column ($B$). After alignment, as shown in the figure, both columns share the same runs [01][0-1], [24][2-4], and [58][5-8]. The aggregation column B is represented with aligned values $[3,3,3]$ for these runs.
  • Step 3: Aggregation. We do this in two steps—(a) perform a parallel aggregation for each segment in the aligned aggregation column by computing $v\times l$ using its values $[3,3,3]$ and run lengths $[2,3,4]$, resulting in $3\times 2,3\times 3,3\times 4]=[6,9,12]$, (b) perform a parallel aggregation for each group-by key by applying torch.scatter on the tensor $[6,9,12]$ using the inverse index $[0,1,0]$. The scatter results in values $6+12=18$ (for group A, index 0) and $9$ (for group B, index 1). We retrieve the group-by keys $[A,B]$ corresponding to the inverse indices $[0,1]$.
A.3. Join Index Generation Example
We illustrate how join indices are generated between columns with different encodings, showing the three-step process of value matching, expansion based on run lengths, and final index generation.
Example 9.
Consider generating Join Indices between a Plain DataColumn and an RLE DataColumn as shown in Figure 15. The Plain DataColumn contains values [A, B, B] at positions 0, 1, and 2. The RLE DataColumn contains values $v=[A,B]$ with start positions $s=[0,2]$, end positions $e=[1,2]$, and run lengths $l=[2,1]$ (l=es+1l=e-s+1).
Appendix B TPC-H Experiment Details
B.1. Query-Specific Column Orderings
Table 7 details the query-specific column orderings used for TPC-H queries to maximize RLE compression effectiveness. Each table is sorted using a global multi-column ordering (equivalent to SQL’s ORDER BY clause) based on the columns involved in each query’s filters and joins.
@@ B.1. Query-Specific Column Orderings
Q1 l_returnflag, l_linestatus, l_shipdate, l_quantity LINEITEM
@@
B.2. TPC-H with Skewed Data
Previous query-specific ordering corresponds to BI scenarios where known queries can be optimized. We also experiment with general ordering using V-order (Microsoft, 2024c) on TPC-H with skew. V-order applies sorting on tables, dictionary encoding, and other techniques to improve compression while supporting fast reads. We experimented with TPC-H using V-order that is not query-specific, on TPC-H SF=50 with skewness parameter z=1. However, only Q1 showed performance speedup, while the rest did not improve. We find that skew alone is insufficient for effective RLE compression. Only Q1 shows speedup because the group-by columns (l_returnflag and l_linestatus) have low cardinality and are naturally RLE-friendly. For other queries, most columns are not suitable for RLE compression, resulting in higher query times due to decompression overhead when operating with Plain columns.
Figure 16. TPC-H query run times with skewed data (z=1) at SF=50, comparing Plain vs Compressed data execution on GPU. Only Q1 shows improvement due to low-cardinality columns suitable for RLE compression.
Appendix C Production Workload Experiment Details
This section provides comprehensive details of our production dataset experiments, including dataset characteristics, compression visualizations, and scalability analysis.
C.1. Fact Table Column Usage
Table 8 shows the columns accessed by each production query evaluated in Section 9.2. The queries operate on a star schema with a fact table containing 2.94 billion rows.
Table 8. Columns of the fact table used by the production queries. Column numbers are anonymized due to proprietary constraints.
@@ C.1. Fact Table Column Usage
Q1 accesses 10 columns and performs 7 semi-joins and 2 PK-FK joins. Q2 and Q3 each access 12 columns and perform 10 semi-joins and 1 PK-FK join. Together, the three queries use 15 distinct columns from the fact table. The total size for these 15 columns is 120.36 GiB with Plain encoding (exceeding the 80 GiB GPU HBM capacity) and 56.84 GiB with compressed encodings.
C.2. RLE Compression Characteristics
The production dataset demonstrates substantial opportunities for RLE compression across multiple columns. Figure 17 shows the detailed compression statistics for the 7 RLE-compressed columns.
@@ C.2. RLE Compression Characteristics
Column 7 exhibits the highest compression with a single run of 2.94B rows (same value for all rows), representing perfect compression. Average run lengths for other columns range from several tens to hundreds of thousands, indicating varying degrees of data locality and repetition.
For example, column 11 has 85.35M runs with an average run length of 34.41. Each run requires 20 bytes for representation—4 bytes for value ($v$), and 8 bytes each for start ($s$) and end ($e$) positions. This totals 1.59 GiB memory for the column, which is significantly less than the 10.94 GiB (= 4 × 2.94B) needed for the Plain representation, achieving a 6.9× compression ratio. Further memory savings are possible by using narrower data types for the start and end positions based on the actual data range.
C.3. Scalability Analysis
To study performance scaling with data size, we study scalability characteristics by varying the input data size. For this we create smaller datasets with the first 5% (0.15B), 20% (0.6B), and 50% (1.5B) rows of the fact table.
Figure 18 shows good performance scalability since GPU query times on compressed data increase much slower with data size compared to GPU query times on plain data. Queries cannot be run on Plain data for the 50% (and 100%) dataset due to running out of GPU HBM capacity, and Q2 and Q3 GPU times for just the 20% dataset exceed that of SQL Server and Analysis Services on CPUs for the 100% dataset! Thus, query execution on compressed data is necessary to unlock the GPU acceleration potential for this workload. We could fit and process even larger datasets in the GPU HBM with our compression schemes.